# The FFT Strikes Back: An Efficient Alternative to Self-Attention

### v3

@@ 4 Experiments
We evaluate our proposed method, FFTNet, comparing it to FNet and standard self-attention-based Transformers. We present results on the Long Range Arena (LRA) benchmark Tay et al. (2021) and the ImageNet classification task. We also provide ablation studies to highlight the contributions of each FFTNet component.
4.1 Training Scheme
We train all models using a combination of standard and advanced data augmentations (e.g., random crops, color jitter, random erasing, etc.) alongside mixup and cutmix strategies for better generalization. The optimization leverages AdamW with a cosine learning-rate schedule and a brief warmup phase. We also apply gradient clipping and optionally employ an exponential moving average (EMA) of parameters. Full details of these hyperparameters and augmentations are available in the accompanying training script.
@@ 4.2 Long Range Arena (LRA) Benchmark
We evaluate on six tasks in LRA: ListOps, Text, Retrieval, Image, Pathfinder, and Path-X. Table 1 reports the accuracy (%) on each task, as well as the average performance across all tasks. Our FFTNet model achieves higher accuracy on most tasks, including a 37.65% accuracy on ListOps (compared to 36.06% for the standard Transformer and 35.33% for FNet). Overall, FFTNet slightly outperforms both baselines on average.
Model ListOps Text Retrieval Image Pathfinder Path-X Avg.
Transformer 36.06 61.54 59.67 41.51 80.38 OOM 55.83
FNet 35.33 65.11 59.61 38.67 77.80 FAIL 55.32
FFTNet (ours) 37.65 66.01 60.21 42.02 80.71 83.25 58.31
Table 1: Accuracy (%) on the Long Range Arena (LRA) benchmark. OOM indicates out-of-memory and FAIL indicates the model could not process the dataset. Our FFTNet obtains the best average accuracy.
@@ 4.3 ImageNet Classification
Next, we evaluate our FFTNetViT variants on the ImageNet classification task, comparing them to standard ViT (self-attention). Table 2 presents the FLOPs and both Top-1 and Top-5 accuracy for each variant (Base, Large, Huge). We omit latency from the table for clarity and show it in a separate figure. Notably, FFTNetViT often achieves lower FLOPs than ViT for comparable model sizes, while maintaining strong accuracy. In most experiments, FNet lags behind both methods on ImageNet, so we focus on comparing our approach to the stronger self-attention baseline.
Variant FFTNetViT ViT
FLOPs Top-1 (%) Top-5 (%) FLOPs Top-1 (%) Top-5 (%)
Base 22.64 79.6 94.9 36.65 79.4 94.8
Large 79.92 82.1 96.2 127.18 81.8 96.0
Huge 166.14 83.2 96.8 261.39 82.9 96.6
Table 2: Comparison of FFTNetViT vs. Standard ViT on ImageNet, grouped by variant. FLOPs are in GFLOPs.
Figure 1: Latency comparison of FFTNetViT vs. Standard ViT for varying batch sizes on ImageNet. FFTNetViT scales faster than standard self-attention.
@@ 4.4 Ablation Studies
We further investigate the contributions of individual FFTNet components by conducting ablation experiments on the ImageNet classification task (Base variant). In particular, we evaluate the following variants:
  • FFTNet (full): The complete model with spectral gating, the adaptive module, and FFT-based filtering.
  • FFTNet without spectral gating: The spectral gating mechanism is removed.
  • FFTNet without adaptive module: The adaptive module is omitted.
  • FFTNet with convolutional replacement: The FFT layer is replaced with a standard convolutional layer.
Table 3 reports the Top-1 accuracy for each variant. As shown, each component contributes positively to the overall performance, with the convolutional replacement yielding the largest degradation.
Variant Top-1 Acc (%) Observation
FFTNet (full) 79.6 Full model with spectral gating, adaptive module, and FFT-based filtering.
   – without spectral gating 78.3 Removing the spectral gating mechanism leads to a noticeable accuracy drop.
   – without adaptive module 77.8 Omitting the adaptive module further reduces performance.
   – FFT replaced with convolution 77.2 Replacing the FFT layer with a convolution results in the largest degradation.
Table 3: Ablation study on the ImageNet classification task (Base variant). Each variant removes or modifies one component of the full FFTNet model.
Overall, these experiments indicate:
  • FFTNet surpasses FNet on both LRA and ImageNet, demonstrating improved accuracy and efficiency.
  • Compared to standard self-attention, FFTNetViT often has lower FLOPs for similar or better performance.
  • Ablations confirm the importance of each FFTNet component (spectral gating, adaptive module).

### v4

@@ 4 Experiments
We present a comprehensive evaluation of our proposed approach, FFTNet, comparing it against both FNet and traditional Transformers based on self-attention. Our experiments cover the Long Range Arena (LRA) benchmark Tay et al. (2021) and ImageNet classification. In addition, we perform ablation studies to highlight the influence of individual components within FFTNet.
4.1 Training Protocol
All models are trained using the AdamW optimizer with a cosine decay schedule, employing an initial warmup phase for the learning rate. We adopt robust data augmentation, including random cropping, color jitter, mixup, and cutmix, and also apply gradient clipping. In certain configurations, we maintain an exponential moving average (EMA) of the model parameters. Detailed hyperparameter choices and augmentation steps can be found in the accompanying training script.
@@ 4.2 Long Range Arena (LRA) Benchmark
We evaluate on six LRA tasks: ListOps, Text, Retrieval, Image, Pathfinder, and Path-X. Accuracy results for each task, as well as the mean accuracy across all tasks, appear in Table 1. We compare the No-Windowing and With Windowing variants of FFTNet against the standard Transformer and FNet. As shown, even the No-Windowing version of FFTNet outperforms both baselines on most tasks. Adding local windowing (STFT) further improves the accuracy slightly.
Model ListOps Text Retrieval Image Pathfinder Path-X Avg. Transformer 36.06 61.54 59.67 41.51 80.38 OOM 55.83 FNet 35.33 65.11 59.61 38.67 77.80 FAIL 55.32 FFTNet (No-Windowing) 37.65 66.01 60.21 42.02 80.71 83.25 58.31 FFTNet (With Windowing) 38.02 66.25 60.64 42.45 80.99 83.64 58.83
Table 1: Accuracy (%) on the Long Range Arena (LRA) benchmark tasks. “OOM” indicates out-of-memory errors; “FAIL” marks cases where models could not process the dataset. FFTNet (With Windowing) achieves the highest average accuracy.
@@ 4.3 ImageNet Classification
We further assess our FFTNetViT variants on ImageNet, comparing them against standard ViT baselines. Table 2 provides FLOPs (in GFLOPs), Top-1 accuracy, and Top-5 accuracy for Base, Large, and Huge model sizes. Latency results (for the No-Windowing variant) are omitted here for brevity but appear separately to illustrate batch-size scalability. In most cases, FFTNetViT uses fewer FLOPs than a similarly sized ViT, while matching or slightly improving accuracy. FNet is less competitive than both FFTNetViT and ViT, so our primary comparison focuses on the latter. As shown, adding local windowing (STFT) brings a small but consistent accuracy boost over the No-Windowing variant.
FFTNetViT (No-Windowing) FFTNetViT (With Windowing) ViT Variant FLOPs Top-1 (%) Top-5 (%) FLOPs Top-1 (%) Top-5 (%) FLOPs Top-1 (%) Top-5 (%) Base 22.64 79.6 $\uparrow 0.2\%$ 94.9 $\uparrow 0.1\%$ 22.64 79.8 $\uparrow 0.4\%$ 95.0 $\uparrow 0.2\%$ 36.65 79.4 94.8 Large 79.92 82.1 $\uparrow 0.3\%$ 96.2 $\uparrow 0.2\%$ 79.92 82.3 $\uparrow 0.5\%$ 96.3 $\uparrow 0.3\%$ 127.18 81.8 96.0 Huge 166.14 83.2 $\uparrow 0.3\%$ 96.8 $\uparrow 0.2\%$ 166.14 83.4 $\uparrow 0.5\%$ 96.9 $\uparrow 0.3\%$ 261.39 82.9 96.6
Table 2: Comparison on ImageNet for Base, Large, and Huge variants. FLOPs are given in GFLOPs. “No-Windowing” vs. “With Windowing” are two configurations of FFTNetViT. The green arrows indicate improvements over the corresponding ViT.
Figure 2: Latency comparison (No-Windowing) of FFTNetViT vs. standard ViT for various batch sizes on ImageNet. FFTNetViT typically exhibits better scaling properties.
Throughput and Speedup Analysis (No-Windowing).
In addition to latency, we compare throughput (images/second) and speedup (ViT latency / FFTNetViT latency) across different batch sizes for Base, Large, and Huge variants. Figure 3 displays these two plots side by side, illustrating:
  • Throughput: FFTNetViT variants consistently process more images per second than standard ViT, and the gap widens at larger batch sizes.
  • Speedup: The ratio of ViT latency to FFTNetViT latency is generally above 1, confirming that FFT-based variants are faster. This advantage becomes more pronounced as the batch size increases.
(a) Throughput (images/second) for FFTNetViT vs. standard ViT.
(b) Speedup (ViT latency / FFTNetViT latency) for FFTNetViT vs. standard ViT.
Figure 3: Comparison of throughput (left) and speedup (right) across various batch sizes (No-Windowing) for Base, Large, and Huge variants of FFTNetViT vs. standard ViT.
@@ 4.4 Ablation Studies
Ablation studies on ImageNet (Base variant, No-Windowing) were conducted to isolate key FFTNet components. Four configurations are examined. The first is the full FFTNet model, which includes spectral gating, the adaptive module, and FFT-based mixing. The second removes spectral gating. The third excludes the adaptive filtering mechanism. The fourth replaces FFT-based mixing with a standard convolution. Table 3 shows the Top-1 accuracy of each variant. All components contribute to performance, with substituting the FFT layer for a convolution causing the largest drop.
Variant Top-1 Acc (%) Observation FFTNet (full) 79.6 Contains spectral gating, adaptive module, and FFT-based filtering. Without spectral gating 78.3 Omitting gating causes a moderate performance decline. Without adaptive module 77.8 Disabling the adaptive module degrades accuracy further. Replace FFT with convolution 77.2 Standard convolution (with a fixed kernel size) leads to the largest drop in performance.
Table 3: Ablation experiments on ImageNet (Base variant, No-Windowing). Each row omits or replaces one key FFTNet component from the full setup.
The results confirm that FFTNet consistently surpasses FNet on both LRA and ImageNet, offering higher accuracy and efficient scaling. When compared to self-attention, FFTNetViT typically has lower FLOPs while delivering comparable or slightly superior accuracy. The ablation experiments highlight the importance of both spectral gating and the adaptive module for peak performance.